Calculate the sample variance and sample standard deviation for the following frequency distribution of heights in centimeters for a sample of 8-year-old boys. If necessary, round to one more decimal place than the largest number of decimal places given in the data. Heights in Centimeters Class Frequency 119.1 - 122.2 46 122.3 - 125.4 38 125.5 - 128.6 24 128.7 - 131.8 19 131.9 - 135.0 27

Heights in Centimeters
Class Frequency
119.1 - 122.2 46
122.3 - 125.4 38
125.5 - 128.6 24
128.7 - 131.8 19
131.9 - 135.0 27
Answer #1

Solution:

Given that,

Class Frequency Midvalue d fd fd2
119.1 - 122.2 46 120.65 -6.4 -294.4 1884.16
122.3 - 125.4 38 123.85 -3.2 -121.4 389.12
125.5 - 128.6 24 127.05 0 0 0
128.7 - 131.8 19 130.25 3.2 60.8 194.56
131.9 - 135.0 27 133.45 6.4 72.8 1105.92
n = 154 \sumfd = -182.4 \sumfd2 = 3573.76

b ) The  sample variance S2

  S2 = (\sum fd2 ) - (( \sum fd )2 / n ) / 1 -n ) * h2

= (3573.76 ( ( -182.4 )2 / 154 ) / 153 * 12

   = (3573.76 - 216.0374 / 153 ) * 1

= (3357.7226 / 153 )* 1

= 21.9459*1

= 21.9459

The  sample variance S2 is 21.9459

b ) The sample standard deviation is S

S = \sqrt{} variance

= \sqrt{} 21.9459

= 4.6846

sample standard deviation is 4.6846